Find the eigenvalues of the given matrices

17. 21] 4t

11. Verify Property 2 for 6 A= 3 -1 2 21 7

Property 2 A matrix is singular if and only if it has a zero eigenvalue
Answer #1

Let A = 2t 4t H - 22 is ginen by Then chiean of A det (A-22=0 X 22 - 2t 4t-) E -> (42-?) + 4t (1=2+) = 0 = 1=24, 27 The Eigen3 6 Ginen 2 =) 7 2 chean of A given by 6 . -X det A-df) to a -1 2-4-1 7-8 using co=C3² 2 262 3 1- - 1 -1-1212.17 7-x-2 2 usinl-d 3 - 1 2- 2-5 5-7 2 3 [A-5) 2(23-02) M o (1–17 [CA-5) (4-2) –(21-5)] =(1-12 [**=+* 7/+10-28+] -3[ 5-31] (1-2) (97-99 +5 )